OpenStudy (anonymous):

need help with answer.. im undecided!! To keep his baseball team hydrated daniel mixes a powdered sports drink w/cold water. For the best teaste the mix should be 96% water. If Daniel has used 16 scoops of powder and added 2 quarts of water, how many gallons of water shoudl he add, given that each scoop is 1/4 cup?

OpenStudy (anonymous):

should he add an additional 5 1/2 gallons of water or 6 gallons of water??

OpenStudy (dangerousjesse):

\(16*.25=4\) cups

OpenStudy (dbzfan836):

that means he should add 96 cups, and you calculate

OpenStudy (dangerousjesse):

\[6.5*.95=x\]

OpenStudy (dbzfan836):

96 cups of water

OpenStudy (dangerousjesse):

Convert \(x\) gallons into \(y\) cups and if \(y=4\), then 6 gallons is right.

OpenStudy (anonymous):

16 scoops x (1/4) cup per scoop = 4 cups powder = 1 quart powder 96% water means water / (water + powder) = .96 let w = quarts of water w / (w + 1) = .96 w = .96(w + 1) w = .96w + .96 .04w = .96 w = 24 quarts = 6 gallons water He already used 2 quarts (1/2 gallon) so he should add 5 1/2 gallons

OpenStudy (anonymous):

I think im going with 5 1/2 gallons

OpenStudy (dbzfan836):

good job, @aballer101

OpenStudy (dangerousjesse):

Yes, that was the right answer :)

OpenStudy (anonymous):

Thank you guys for the help!